# Radiopurity screening of materials for rare event searches by neutron activation at the TRIGA reactor of Pavia

### v1

@@ 3.5 Expected NAA sensitivity at Milano-Bicocca
[…]
Finally, we estimate the sensitivities regarding the concentration of natural contaminants that could potentially be achieved with our experimental setup in the absence of interfering activation products. It is important to note that this estimate serves primarily for illustrative purposes and has a key limitation: it does not account for correlated background, which varies depending on the sample under investigation. For this calculation, we start from Eq. 4 and we explicate $R=\mathcal{N}\,\sigma_{\text{eff}}\,\Phi$. We assume a 6-hour irradiation in the Lazy Susan facility at the TRIGA reactor of Pavia and we use the neutron flux and effective cross section values obtained from the measurements presented in Sect. 4 (see Table 5, 2nd column, and Table 6). Furthermore, we set the number of counts in the selected $\gamma$-peak equal to a $3\sigma$ Poissonian statistical fluctuation of the number of background counts, ensuring clear identification of the activation product within the sample.
In Fig. 1 we show the attainable sensitivity for the concentration of natural contaminants as a function of measuring time with the BEGe detector, assuming a generic plastic sample weighing 10 g. For illustration purposes, we considered four different cooling times, ranging from a minimum of 4 hours to a maximum of 1.5 times the half-life of the activated product under study, and measuring times up to 3 times the same half-life. As one can see, concentrations ranging from sub-ppt (for 40K) to ppt (for 232Th) level concentrations seem achievable with our experimental setup with a $3\sigma$ confidence level.
@@ 4.1 Experiment description
Figure 2: (a) Al-Co and Al-Au samples sandwiched between two acrylic cylinders (label no.1) and positioned above an empty vial (label no.2). (b) Stack of 8 couples of Al-Co and Al-Au samples positioned inside small polyethylene vials.
@@
Table 3: List of Al-Co (Co 0.1% wt.) and Al-Au (Au 0.134% wt.) samples irradiated in the Lazy Susan (LS) facility of the TRIGA Mark II reactor in Pavia. The experimental uncertainty of Al-Co mass is ±$\pm 0.01$ mg, whereas that of Al-Au samples is ±$\pm 0.003$ mg. In the last column we report the position of samples measured from the bottom of the holder with ±$\pm 0.1$ cm precision.
@@ 4.2 Activation rate results
197Au (n,$\gamma$) 198Au $(5.679\pm 0.009)\times 10^{14}$ $(5.167\pm 0.009)\times 10^{14}$ $9.9\pm 0.2$
59Co (n,$\gamma$) 60Co $(5.018\pm 0.029)\times 10^{14}$ $(4.493\pm 0.024)\times 10^{14}$ $11.7\pm 0.8$
27Al (n,$\alpha$) 24Na $(2.973\pm 0.011)\times 10^{9}$ $(2.889\pm 0.011)\times 10^{9}$ $2.9\pm 0.5$
@@
First of all, we compare the results obtained for $R$ from the measurements on the two different HPGe detectors. The differences are within ±$\pm 3\%$ for all samples, pointing out that our MC tool for $\varepsilon_{\gamma}$ estimation can reach the same accuracy level certified for the activity of the multi-$\gamma$ sources used in our laboratory for HPGe calibration.
@@
In Table 4, we report our best estimates [^4] Computed as the weighted average of the $R/m$ results from the two HPGe measurements. for the SSA of samples no.1 and no.2 irradiated with and without a surrounding moderator, respectively. In the last column, we highlight that the (n,$\gamma$) reactions are characterized by activation rates $\sim 10\%$ higher when the samples are surrounded by a moderator, whereas the difference is lower for the (n,$\alpha$) reaction induced by fast neutrons. Since the samples no.1 and no.2 have been irradiated in two adjacent channels of the Lazy Susan facility, we have considered the hypothesis that there is non-uniformity of flux between the channels. However, given the simulation results shown later in Section 4.4, we can expect flux variations of $\mathcal{O}(1\%)$ between channels LS–27 and LS–28. Therefore, we can state the observed $\sim 10\%$ differences for (n,$\gamma$) activation rates are due to the effect of the moderator. In particular, they can be interpreted as the consequence of a local flux enhancement due to elastic scattering in the moderator, which causes more neutrons, especially thermal ones, to pass back and forth through the sample.
@@ 4.3 Neutron flux results
Figure 5: (a) Neutron flux spectra, shown in lethargic scale and normalized to 1, obtained from the MCNP simulation model of the TRIGA Mark II reactor in Pavia at various distances from the center of the core along the z-axis in the Central Channel (CC), Rabbit (Rab), and Lazy Susan (LS) irradiation facilities, and in 3 annulus volumes in the reflector region between the core and the Lazy Susan (for a description of the TRIGA reactor geometry refer to AbsoluteFlux ). All the other spectra (not shown here) belonging to the set of guess spectra used in the unfolding analysis are characterized by similar shapes, comprised between the most and least thermalized spectra shown here. (b) Spectra, normalized to 1, obtained at the end of the unfolding analysis of the 10 samples irradiated in different positions and conditions in the LS facility described in this paper.
@@
In order to identify, among spectra produced with the MCNP simulations, the one that better fits the unfolding results, we implement a $\chi^{2}$ minimization and we apply the iterative procedure described in Ref. Chiesa:2019buj . After a few iterations ($<10$), the unfolding solution stabilizes at values consistent with one of the spectra available in the MCNP set.
@@
Conversely, we find appreciable differences when the neutron flux results are considered on an absolute scale. Indeed, as shown in Table 5, a relative difference of $(12\pm 3)\%$ is obtained for the total flux measured with or without the moderator. This difference arises from a variation of the thermal and epithermal flux, whereas the fast flux is compatible in the two measurements, as expected. We point out that the thermal and epithermal fluxes are highly anti-correlated (the correlation coefficient is $-0.7$), so their sum can be determined with a smaller uncertainty ($\sim 2.5\%$) than each of them taken separately.
@@ 4.4 Discussion
  • thermal: $64\%-67\%$ (±$\pm 3\%$)
  • epithermal: $24\%-27\%$ (±$\pm 3\%$)
  • fast: $7.6\%-10\%$ (±$\pm 0.4\%$)
@@
A direct application of the neutron flux characterization conducted with this methodology, that combines the accuracy of the experimental measurements with the precision of the Monte Carlo simulations in determining the intra-group spectral shapes, is the calculation of the effective cross sections and the activation rates of any nuclide of interest. We present an example of such application for the most relevant nuclides in the framework of radiopurity assays measurements. We use all the unfolded spectra reported in Fig. 5(b) to calculate the effective cross section for (n,$\gamma$) reactions on 41K, 232Th and 238U. By calculating the effective cross sections with the different spectra, it turns out that they are compatible with each other within their uncertainties. This outcome is easily understandable given the similarity of the unfolded spectra. In Table 6 we report the average values of the effective cross sections.
@@ 5 Conclusions
The sensitivities on 40K, 232Th and 238U contaminations that we can potentially achieve in the absence of interfering activation products is $<0.01$ ppt for 40K, $<1$ ppt for 238U and of the order of a few ppt for 232Th contaminations. Based on our experience, it is quite common to activate interfering nuclides that worsen the sensitivity in the search for natural radionuclides. Therefore, when possible, we employ radiochemical sample treatments and background reduction techniques based on $\beta-\gamma$ and $\gamma-\gamma$ coincidence measurements, which can further improve our sensitivity and will be the subject of future articles.

### v2

@@ 3.5 Expected NAA sensitivity at Milano-Bicocca
Finally, we estimate the sensitivities for 40K, 238U, and 232Th, i.e. the smallest concentrations of natural contaminants that could potentially be detected with our experimental setup in the absence of interfering activation products. It is important to note that this estimate serves primarily for illustrative purposes and has a key limitation: it does not account for correlated background, which varies depending on the sample under investigation. For this calculation, we combine Eq. 4, 12 and 13 to determine the activation rate $R^{*}$ corresponding to the Currie’s detection limit (see Sect. 2.1), which ensures a clear identification of the $\gamma$-peak of the activation product.
$R^{*}=\dfrac{C^{*}}{\varepsilon_{\gamma}\,I_{\gamma}\,f(\Delta E)}\dfrac{\lambda}{(1-e^{-\lambda t_{\text{irr}}})e^{-\lambda t_{\text{c}}}(1-e^{-\lambda t_{\text{m}}})}$ (14)
Then, we use $R^{*}$ to calculate the corresponding number of target isotopes $\mathcal{N}^{*}=R^{*}/(\sigma_{\text{eff}}\,\Phi)$ and, from these, the mass of the contaminant, which is finally divided by the mass of the sample to get the contaminant concentration.
In Fig. 1 we show the attainable sensitivity for the concentration of natural contaminants as a function of measuring time with the BEGe detector, considering a generic plastic sample with $m_{sample}=10$ g and assuming a 6-hour irradiation in the Lazy Susan facility at the TRIGA reactor of Pavia. To this aim, we use the neutron flux and effective cross section values obtained from the measurements presented in Sect. 4 (see Table 5, 2nd column, and Table 6). For illustration purposes, we considered four different cooling times, ranging from a minimum of 4 hours to a maximum of 1.5 times the half-life of the activated product under study, and measuring times up to 3 times the same half-life.
As one can see, concentrations ranging from a few ppq (for 40K) to a few ppt (for 232Th) are detectable with our experimental setup if the HPGe background is not increased by other activation products.
@@ 4.1 Experiment description
Figure 2: (a) Al-Co and Al-Au samples sandwiched between two acrylic cylinders (label no.1) and positioned above an empty vial (label no.2). (b) Stack of 8 couples of Al-Co and Al-Au samples positioned inside small polyethylene vials. (c) Scheme (not to scale) of the setup in the top left photo. (d) Scheme (not to scale) of the setup in the top right photo.
@@
Table 3: List of Al-Co (Co 0.1% wt.) and Al-Au (Au 0.134% wt.) samples irradiated in the Lazy Susan (LS) facility of the TRIGA Mark II reactor in Pavia. The experimental uncertainty of Al-Co mass is $\pm 0.01$ mg, whereas that of Al-Au samples is $\pm 0.003$ mg. In the last column we report the position of samples measured from the bottom of the holder with $\pm 0.1$ cm precision.
@@ 4.2 Activation rate results
[…]
@@
First of all, we compare the results obtained for $R$ from the measurements on the two different HPGe detectors. The differences are within $\pm 3\%$ for all samples, pointing out that our MC tool for $\varepsilon_{\gamma}$ estimation can reach the same accuracy level certified for the activity of the multi-$\gamma$ sources used in our laboratory for HPGe calibration.
@@
[…]
@@ 4.3 Neutron flux results
Figure 5: (a) Neutron flux spectra, shown in lethargic scale and normalized to 1, obtained from the MCNP simulation model of the TRIGA Mark II reactor in Pavia at various distances from the center of the core along the z-axis in the Central Channel ($C$), Rabbit (Rab), and Lazy Susan (LS) irradiation facilities, and in 3 annulus volumes in the reflector region between the core and the Lazy Susan (for a description of the TRIGA reactor geometry refer to AbsoluteFlux ). All the other spectra (not shown here) belonging to the set of guess spectra used in the unfolding analysis are characterized by similar shapes, comprised between the most and least thermalized spectra shown here. (b) Spectra, normalized to 1, obtained at the end of the unfolding analysis of the 10 samples irradiated in different positions and conditions in the LS facility described in this paper.
@@
In order to identify, among spectra produced with the MCNP simulations, the one that better fits the unfolding results, we implement a $\chi^{2}$ minimization and we apply the iterative procedure described in Ref. Chiesa:2019buj . After a few iterations (<10<10), the unfolding solution stabilizes at values consistent with one of the spectra available in the MCNP set.
@@
Conversely, we find appreciable differences when the neutron flux results are considered on an absolute scale. Indeed, as shown in Table 5, a relative difference of $(12\pm 3)\%$ is obtained for the total flux measured with or without the moderator. This difference arises from a variation of the thermal and epithermal flux, whereas the fast flux is compatible in the two measurements, as expected. We point out that the thermal and epithermal fluxes are highly anti-correlated (the correlation coefficient is 0.7-0.7), so their sum can be determined with a smaller uncertainty ($\sim 2.5\%$) than each of them taken separately.
@@ 4.4 Discussion
  • thermal: $64\%-67\%$ ($\pm 3\%$)
  • epithermal: $24\%-27\%$ ($\pm 3\%$)
  • fast: $7.6\%-10\%$ ($\pm 0.4\%$)
@@
[…]
@@ 5 Conclusions
The sensitivities on 40K, 232Th and 238U contaminations that we can potentially achieve in the absence of interfering activation products is <0.01<0.01 ppt for 40K (<0.1<0.1 ppb for elemental K), <1<1 ppt for 238U and of the order of a few ppt for 232Th contaminations. Based on our experience, it is quite common to activate interfering nuclides that worsen the sensitivity in the search for natural radionuclides. Therefore, when possible, we employ radiochemical sample treatments and background reduction techniques based on $\beta-\gamma$ and $\gamma-\gamma$ coincidence measurements, which can further improve our sensitivity and will be the subject of future articles.
@@ Acknowledgments
Data availability statement
The authors declare that the data supporting the findings of this study are available within the paper. The numerical data used to produce the figures are available from the corresponding author on reasonable request.